# Aux-Think: Exploring Reasoning Strategies for Data-Efficient Vision-Language Navigation

### v2

@@ Abstract
Vision-Language Navigation (VLN) is a critical task for developing embodied agents that can follow natural language instructions to navigate in complex real-world environments. Recent advances in VLN by large pretrained models have significantly improved generalization and instruction grounding compared to traditional approaches. However, the role of reasoning strategies in navigation—an action-centric, long-horizon task—remains underexplored, despite Chain-of-Thought (CoT) reasoning’s demonstrated success in static tasks like visual question answering. To address this gap, we conduct the first systematic evaluation of reasoning strategies for VLN, including No-Think (direct action prediction), Pre-Think (reason before action), and Post-Think (reason after action). Surprisingly, our findings reveal the Inference-time Reasoning Collaps issue, where inference-time reasoning degrades navigation accuracy, highlighting the challenges of integrating reasoning into VLN. Based on this insight, we propose Aux-Think, a framework that trains models to internalize structured reasoning patterns through CoT supervision, while inferring action directly without reasoning in online prediction. To support this framework, we release R2R-CoT-320k, the first Chain-of-Thought annotated dataset for VLN. Extensive experiments show that Aux-Think reduces training effort greatly and achieves the best performance under the same data scale.
(a) Comparison of navigation performance of different reasoning strategies, which are only trained on R2R-CoT-320k. The proposed Aux-Think (Ours) method consistently outperforms other reasoning strategies.
@@ 1 Introduction
To address the IRC issue, we propose Aux-Think, inspired by the human learning Dual-process evans2003two : during training, humans often rely on explicit reasoning to understand principles, but during execution, they focus on actions without consciously recalling those principles, much like a driver who no longer recites traffic rules while driving. Similarly, Aux-Think uses CoT as an auxiliary signal during training to guide the model in internalizing reasoning patterns. At inference time, the model no longer generates explicit reasoning, but instead directly predicts actions based on the internalized reasoning learned during training. This separation between learning and execution improves decision focus, reduces inference overhead and hallucinations, and leads to more accurate and stable navigation.
@@
  • New Method: We propose Aux-Think, a novel training paradigm that uses CoT as auxiliary supervision while maintaining No-Think inference, achieving superior performance over other reasoning methods. Aux-Think pioneers a new perspective on CoT utilization.
@@ 3.4 Aux-Think: Reasoning-Aware Co-Training Strategy
$L^{Act}_{\tau}(\theta)=-\sum_{t=0}^{T}\sum_{k=0}^{n}\log\pi_{\theta}(a_{t+n}^{*}|\mathcal{O}_{t},o_{t},I_{\tau})$ (4)
@@ 4.3 Comparison on VLN-CE Benchmarks
In large model-based methods, we additionally mark the amount of data used by the method in addition to the R2R-CE training split (Extra Data). We further scale up by the RxR training split (600K), and our performance remains comparable to those using a much larger amount of training data, whose training data includes general visual data and real-world data from the Internet.
Table 1: Comparison of different methods on the R2R Val-Unseen split. Observations used include Monocular (Mono.) and Panoramic view (Pano.). $\ast$ indicates methods based on the waypoint predictor hong2022bridging . $\dagger$ indicates methods without using LLMs. indicates the models using the training data only from R2R-CE training split, and we compare with the results reported in their paper for a fair evaluation. indicates using extra DAgger ross2011reduction and instruction reasoning data but without RxR data, which is the same setting with NaVid zhang2024navid . The training data structures of traditional non-LLM-based methods are quite different, so we do not compare them with their extra data.
@@
NaVidzhang2024navid RSS2024 $\checkmark$ 5.47 49.1 37.4 35.9 180K
Aux-Think (ours) - $\checkmark$ 6.02 53.0 46.1 41.5 180K
@@
Aux-Think (ours) - $\checkmark$ 5.88 54.9 49.7 41.7 600K
@@
To further assess the ability of Aux-Think, we evaluate it on the RxR-CE ku2020room Val-Unseen split (Table 2). Compared to R2R-CE, RxR-CE includes more natural instructions and longer trajectories, making it a more realistic and challenging benchmark. Aux-Think achieves strong overall performance, particularly on the OSR metric, where it surpasses Uni-NaVid while using much fewer training data (920K vs. 5900K). This demonstrates the effectiveness of reasoning supervision under limited data. It is noted that performance on NE and SPL is relatively modest. This is likely due to the model’s internalized reasoning behavior, learned during CoT training, which encourages broader exploration. Although this can lead to longer paths and reduced efficiency, it improves SR and OSR by increasing the likelihood of reaching the goal, especially in unfamiliar environments.
Table 2: Comparison of different methods on the RxR Val-Unseen split. $\dagger$ indicates methods without using LLMs. The training data structures of traditional non-LLM-based methods are quite different, so we do not include their training data.
@@
Aux-Think (ours) - $\checkmark$ 6.91 60.1 49.3 37.9 920K

### v3

@@ Abstract
Vision-Language Navigation (VLN) is a critical task for developing embodied agents that can follow natural language instructions to navigate in complex real-world environments. Recent advances in VLN by large pretrained models have significantly improved generalization and instruction grounding compared to traditional approaches. However, the role of reasoning strategies in navigation—an action-centric, long-horizon task—remains underexplored, despite Chain-of-Thought (CoT) reasoning’s demonstrated success in static tasks like visual question answering. To address this gap, we conduct the first systematic evaluation of reasoning strategies for VLN, including No-Think (direct action prediction), Pre-Think (reason before action), and Post-Think (reason after action). Surprisingly, our findings reveal the Inference-time Reasoning Collaps issue, where inference-time reasoning degrades navigation accuracy, highlighting the challenges of integrating reasoning into VLN. Based on this insight, we propose Aux-Think, a framework that trains models to internalize structured reasoning patterns through CoT supervision, while inferring action directly without reasoning in online prediction. To support this framework, we release R2R-CoT-320k, the first Chain-of-Thought annotated dataset for VLN. Extensive experiments show that Aux-Think reduces training effort greatly and achieves the best performance on the navigation success rate.
(a) Comparison of navigation performance of different reasoning strategies, which are only trained on R2R-CoT-320k without receding-horizon action planning. The proposed Aux-Think (Ours) method consistently outperforms other reasoning strategies.
@@ 1 Introduction
To address the IRC issue, we propose Aux-Think, inspired by the dual-process theory of human learning evans2003two : during training, humans often rely on explicit reasoning to understand principles, but during execution, they focus on actions without consciously recalling those principles, much like a driver who no longer recites traffic rules while driving. Similarly, Aux-Think uses CoT as an auxiliary signal during training to guide the model in internalizing reasoning patterns. At inference time, the model no longer generates explicit reasoning, but instead directly predicts actions based on the internalized reasoning learned during training. This separation between learning and execution improves decision focus, reduces inference overhead and hallucinations, and leads to more accurate and stable navigation.
@@
  • New Method: We propose Aux-Think, a novel training paradigm that uses CoT as auxiliary supervision while maintaining No-Think inference, achieving superior performance over other reasoning methods. Aux-Think pioneers a new perspective on CoT utilization and achieves the best performance on the navigation success rate.
@@ 3.4 Aux-Think: Reasoning-Aware Co-Training Strategy
$L^{Act}_{\tau}(\theta)=-\sum_{t=0}^{T}\sum_{k=0}^{n}\log\pi_{\theta}(a_{t+k}^{*}|\mathcal{O}_{t},o_{t},I_{\tau})$ (4)
@@ 4.3 Comparison on VLN-CE Benchmarks
In large model-based methods, we additionally mark the amount of data used by the method in addition to the R2R-CE training split (Extra Data). We further scale up by the RxR training split (600K), DAgger data (500K) and web data (500K), and our performance achieves the SOTA Success Rate (SR) to those using a much larger amount of training data.
Table 1: Comparison of different methods on the R2R Val-Unseen split. Observations used include Monocular (Mono.) and Panoramic view (Pano.). $\ast$ indicates methods based on the waypoint predictor hong2022bridging . $\dagger$ indicates methods without using LLMs. indicates the models using the training data only from R2R-CE training split, and we compare with the results reported in their paper for a fair evaluation. The training data structures of traditional non-LLM-based methods are quite different, so we do not compare them with their extra data.
@@
Aux-Think (ours) - $\checkmark$ 6.08 60.0 54.8 46.9 1600K
@@
To further assess the ability of Aux-Think, we evaluate it on the RxR-CE ku2020room Val-Unseen split (Table 2). Compared to R2R-CE, RxR-CE includes more natural instructions and longer trajectories, making it a more realistic and challenging benchmark. Aux-Think achieves strong overall performance, particularly on the Success Rate (SR) metric, where it surpasses Uni-NaVid and NaVILA while using much fewer training data (1920K vs. 5900K and 3100K). This demonstrates the effectiveness of reasoning supervision under limited data. It is noted that performance on NE and SPL is relatively modest. This is likely due to the model’s internalized reasoning behavior, learned during CoT training, which encourages broader exploration. Although this can lead to longer paths and reduced efficiency, it improves SR and OSR by increasing the likelihood of reaching the goal, especially in unfamiliar environments.
Table 2: Comparison of different methods on the RxR Val-Unseen split. $\dagger$ indicates methods without using LLMs. The training data structures of traditional non-LLM-based methods are quite different, so we do not include their training data in this table.
@@
Aux-Think (ours) - $\checkmark$ 6.24 61.9 52.2 40.2 1920K
@@ A.2 R2R-CoT-320k
To provide a deeper quantitative understanding of our proposed R2R-CoT-320k dataset, we present statistics on CoT content and complexity. As shown in Fig. 5(a), the word cloud reveals frequent reasoning patterns grounded in spatial semantics, such as “doorway,” “current observation,” “hallway,” “turning,” and “goal.” These tokens suggest that the dataset captures rich, step-by-step reasoning tightly aligned with embodied navigation semantics.
Fig. 5(b) shows the distribution of CoT lengths, where most reasoning chains fall within the 200–300 word range, but with a long tail reaching beyond 450 words. This indicates that the dataset covers both concise and highly detailed reasoning processes, posing a greater challenge than typical short-form CoT datasets used in static tasks.
Overall, R2R-CoT-320k represents the first large-scale reasoning-augmented dataset for VLN with diverse, high-coverage CoT annotations. It offers a valuable benchmark to study the role of language-based reasoning in long-horizon, partially observable navigation tasks.
Figure 5: (a) Word cloud of Chain-of-Thought in the R2R-CoT-320k dataset, highlighting frequent visual and spatial reasoning patterns. (b) Distribution of CoT lengths (in word count), showing a wide and diverse range of reasoning complexity.